If the energy stored in a 1/8 F capacitor is 25 J, find the voltage and charge.


1
Expert's answer
2021-12-07T10:37:26-0500

energy:

"E=CU^2\/2"


voltage:

"U=\sqrt{2E/C}=\sqrt{2\cdot25\cdot8}=20" V


charge:

"q=CU=20\/8=2.5" C